A doctor can see 10 patients in 2.5 hours.
A. How many patients can be seen in an hour? I know this is 4
B. Write an equation to represent the relationship between patients and hours.

Answer by Nate(3500)

A.) 10/2.5 = 100/25 = 4 patients
B.) f(x) = 4x where f(x) is the amount of patients in x amount of hours